A source of power 600 kW emits photon which incident on a surface 2.5 m away out of which 50% is reflected back, what is the radiation pressure on the surface?

Grade:12th pass

1 Answers

Susmita
5 years ago
Intensity at 2.5 m way is
I=600/(2.52)=96kW/m2
Pressure p=2I/c where c is speed of light in free space=3×108m/s
p=6.4×10^-4 N/m2
please approve if helped.